TIME-RESOLVED ROTATIONAL ENERGY TRANSFER AND SPECTRAL LINE BROADENING IN ACETYLENE. A HIGH RESOLUTION RAMAN STUDY.

\underline{J. L DOMENECH}, R. Z. MARTINEZ, D. BERMEJO, Instituto de Estructura de la Materia (CSIC), Dept. of Molecular Pysics, Serrano 123, E-28006 Madrid, Spain.

The quasi-cw stimulated Raman technique is a powerful tool for the determination of collisional broadening and linemixing parameters, which bear a very direct relationship with the rotational energy transfer rate constants matrix at the state-to-state level. Indeed, the broadening coefficients and the analysis of line-mixing profiles, frequently combined with a rate-law analysis, have provided a wealth of information on collision physics for many gas systems throughout the years. We report on new high resolution measurements of the broadening coefficients on the Q-branch of the ν_{2} band of acetylene.

Furthermore, last year at this meeting, we reported on a new technique for the determination of state-to-state rotational energy transfer rate constants. based on a time-resolved double-resonance Raman-Raman scheme. It uses the quasicontinuous stimulated Raman-loss technique for the probe stage. Through a systematic recording of the time evoultion of rotational populations from a single initially pupulated J level, it allows the direct determination of most elements of the rate-constant matrix without resorting to fitting or scaling laws.

The comparison of the results obtained from from the high resolution spectroscopic approach with those from the timeresolved dynamics experiment provides insight into the collision mechanisms and some of the assumptions or simplifications usually made in rate-law analysis.